November 2, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, November 2, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates - mcsmear 20160323
  3. Data Validation
  4. GlueX Detector Performance Document

Upcoming Calibrations

  • Add timing information to FCAL clustering/inner ring calibrations? (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • TAGH alignment (Nathan)
  • Tagger energy map (Richard J.?)
  • TOF alignment (Beni/Simon)


  1. Announcements
  2. Reconstruction Launch Readiness
  3. Run Planning
  4. Calibration Updates
  5. Simulations
    • Data/Simulation comparisons
  6. AOB


Attending: Sean (NU); Simon, Adesh, Alex A., Eugene, Elton, Thomas, Mike S., Lubomir, Mark I., Beni, Alex B., Mark D. (JLab); Will M., Curtis (CMU); Matt S. (IU); Nacer (GSI); Cris (MIT); Justin (W&M)

  1. Announcements
    • Beam is expected in ~2 weeks
  2. Reconstruction Launch Readiness
    • Alex A. reported that the last week was busy with launch preparations. Updates included new FDC time calibrations, TAGM timing updates, and fixes to the CDC geometry.
    • Monitoring version 9 has been started, should be finished within a day. If the data is promptly looked at, we can start the next reconstruction run this week.
    • Mark I.: A new official version of HDDS should be tagged.
    • AA: tried doing timing alignment to individual FDC planes, saw some improvement but needs more study
    • Mike S.: New FDC calibration procedures are solid, will check in new scripts soon.
  3. Run Planning
    • Sean, Naomi and David have ironed out issues with filling out the BOR record with the new firmware versions. This data is properly processed in sim-recon, but David is still ironing out some issues about making it available.
      • Mike S.: Is this information used for production data? Sean: Yes, it's needed to get the number of samples used for calculation the pulse integrals and the pedestals to properly perform the pedestal subtraction.
      • Mike S.: Isn't there other configuration information in each event? Sean: Yes, but is inconsistent with BOR information. Needs to ask David about this, but it is a low priority - we should rely on the BOR data which reflects the configurations actually set in the firmware in any case.
    • Several people have been looking at the new f250 firmware output. Sean has updated the emulation code in sim-recon. Looking at a file of BCAL cosmic data, he sees consistent results for pulse parameters, except the timing is wrong at the sub-percent level. The status flags make no sense, though. Alex S. is aware of these issues, and had found a couple of particular bugs, and is waiting for a new firmware version.
    • Naomi has been looking at the new f125 firmware, but was not able to be at the meeting. She has said that things are looking better. Beni will check to make sure that all of the firmware is up to date.
    • Sean has also been working on testing the new version of the calibration train processing with runs from the spring. He is also adding a pass over the skimmed PS data. He is currently working on resolving some CCDB bugs with Dmitry, and hopes to show something next week.
  4. Calibration Updates
    • FCAL - Adesh updated the timing offsets before the last monitoring launch, and improvements in timing resolution have been seen. In updating the monitoring pages, Thomas noted that he saw some runs where the FCAL was recording a large number of presumably noise hits, which Adesh said was after a DAC threshold calibration. Matt S.: One should look at the clusterizer output to see if this affects physics analysis, not the hit level. Sean: Send me a list of runs, will add it to the "known issues" list.
    • BCAL - Mark Dalton has updated the per-channel timing calibrations. He is working on improving the time resolution towards the design goal, but sees several challenges. He will report on these in detail in tomorrow's calorimetry meeting.
    • CDC - Mike determined updated values for the location of the CDC using Lubomir's method of projecting reconstructed tracks in the FDC to the CDC, but using a different track fitting program. The results look better.
        • Matt S.: Noticed that the event vertex after kinematic fit is shifted by ~1 mm. Mike S.: This is a known problem. Still working on how to perform inter-detector alignment.
    • TOF - Beni tested the new non-linear correction calibration table using one board overnight. The results look very good, he's now updating all of the boards. Then Paul M. can calibrate the RF system.
    • Start Counter - Sean will follow up with Mahmoud on the status of the efficiency calculation, with a cc to Mark Ito.
    • TAGM - Alex B. has updated the calibrations in CCDB, and is ready to go for the next run. Elton & Mark D. asked about the TAGM time walk procedure, which is done relative to the RF time rather than the fADC time. They are interested in knowing the time walk of the fADC for BCAL calibrations. Alex replied that this method was used to deal with low amplitude pulses, and will take a look to see if he can get the information requested.
  5. Simulations
    • BCAL - Will M. reported that is looks like all of the bugs have been fixed and he can now match data and simulation without changing any parameters to apparently unphysical values. Sean asked him to show some comparisons, perhaps at tomorrow's calorimetry meeting.
    • FCAL - Adesh has been looking at simulations, and finds a pi0 peak with width ~6 MeV compared to ~8 MeV in data. He needs more statistics to do a more detailed analysis. The timing resolution also needs to be updated, and he plans to do this after the next monitoring launch. Sean said that Adesh should provide the number of events needed, and it might be time to think about another reconstruction run.
    • After some brief discussion, Sean suggested that after we start taking steady production data, we should come up with a plan to update the tracking in simulations.